# DiLA: Enhancing LLM Tool Learning with Differential Logic Layer

### v3

@@ Appendix B Details of the Initialization
After parsing logic formulas from the problem description, we can directly ask LLM to generate a possible solution according to its language understanding. The possible question prompt can be “What is the logical solution based on the premises?” or “Can you provide a valid answer from these given constraints?”. Here we show a prompt example for a simple boolean satisfiability problem in Table 4.
Example prompt of initial solution generation
Question: Consider three Boolean variables: A, B, and C. The following constraints must be satisfied: 1. At least one of A or B is true; 2. At least one of B or C is true; 3. A and C cannot both be true; 4. B cannot be true if A is true. Can you analyze this Boolean satisfiability problem and give a valid answer?
LLM response: Let’s break down the constraints:
1. At least one of A or B is true: $A\vee B$.
2. At least one of B or C is true: $B\vee C$.
3. A and C cannot both be true: $\neg(A\wedge C)=\neg A\vee\neg C$.
4. B cannot be true if A is true: $\neg(A\wedge B)=\neg A\vee\neg B$.
Therefore, the overall constraint formula would be $(A\vee B)\wedge(B\vee C)\wedge(\neg A\vee\neg C)\wedge(\neg A\vee\neg B).$
Now, let’s analyze these constraints:
From constraint 4, we know that if A is true, B must be false. This means that if A is true, the only way to satisfy constraint 1 is to have B false and C true (since at least one of B or C must be true).
From constraint 3, we know that A and C cannot both be true. This means that if A is true, C must be false.
So, we have two possible scenarios:
Scenario 1: A is true, B is false, and C is false.
Scenario 2: A is false, B is true, and C is true.
Let’s check if these scenarios satisfy all the constraints:
Scenario 1:
1. Constraint 1: $A\vee B$ is true (A is true).
2. Constraint 2: $B\vee C$ is false (both B and C are false), but this is okay because we can still satisfy it by making C true.
3. Constraint 3: $\neg A\vee\neg C$ is true (A is true, C is false).
4. Constraint 4: $\neg A\vee\neg B$ is true (A is true, B is false).
Scenario 2:
1. Constraint 1: $A\vee B$ is true (B is true).
2. Constraint 2: $B\vee C$ is true (B is true).
3. Constraint 3: $\neg A\vee\neg C$ is true (A is false, C is true).
4. Constraint 4: $\neg A\vee\neg B$ is true (A is false, B is true).
Both scenarios satisfy all the constraints. Therefore, the Boolean satisfiability problem has two valid solutions:
1. A = true, B = false, C = false
2. A = false, B = true, C = true
These are the two possible assignments of values to A, B, and C that satisfy all the constraints.
Table 4: Example of generating an initial solution for a boolean satisfiability problem.
From this prompt response by the LLM, we have two key observations. First, the LLM can successfully understand semantic constraints, transform them into logical formulas, and perform logical reasoning. The possible solution, i.e., variable assignment, provided by the backbone LLM is reasonable and can satisfy some of the given constraints based on its language understanding. However, LLMs still make some mistakes when performing logical reasoning. For instance, in Scenario 1, the solution provided does not satisfy all constraints, despite the LLM’s assertion that it does (as indicated by the words in red in the response). Therefore, it is necessary to design a powerful tool to check the validity of the given solution and further refine it until all constraints are satisfied.
Appendix C Dataset Generation
@@
Graph coloring benchmarks belong to the class of NP-complete problems with which exact heuristic-based solvers struggle. To test the effectiveness of our proposed DiLA, we built our dataset using GrinPy [^6] https://pypi.org/project/grinpy/ for common graph operations. Once a successful candidate is found, it is compiled into the standard DIMACS format and appended with a comment containing its precalculated chromatic number. Specifically, before inputting to LLMs, we transform the graph connections in DIMACS format (e.g., 0 3) into natural language statements (e.g., Vertex 0 is connected to vertex 3.). For the experiments in Section 4.2, we generated 100 graph coloring instances with node counts ranging from 10 to 200 and edge counts from 30 to 480. Notably, all these artificial graphs can be successfully colored with 3 colors, providing an ideal and easily solvable setting. An example of one of the graphs with 10 nodes and 30 edges as well as its 3-coloring solution is shown in Figure 5.
@@
Table 5: Edge connections in language for a 3-coloring problem.

### v4

@@ Appendix A More Evaluations on Natural Language Reasoning
After parsing logic formulas from the problem description, we can directly ask LLM to generate a possible solution according to its language understanding. The possible question prompt can be “What is the logical solution based on the premises?” or “Can you provide a valid answer from these given constraints?”.
Appendix B Discussions on UNSAT
In DiLA, we indeed find the solution for the MaxSAT problem in Equation 1, and we denote the maximum satisfiable set of clauses as $\phi$. For an UNSAT problem, the complementary set of $\phi$, i.e., the remaining unsatisfied clauses $\bar{\phi}$, must include at least one clause from minimally unsatisfiable subformulas, i.e., UNSAT core. Therefore, based on the MaxSAT results from DiLA, we can design a procedure to iteratively find the UNSAT core:
  1. 1.
    MaxSAT Solution: DiLA first finds the maximal satisfiable subset of clauses $\phi$ for the problem.
  2. 2.
    Source of Conflict: The complementary unsatisfied clauses $\bar{\phi}$ must contain at least one clause from UNSAT cores.
  3. 3.
    Iterative UNSAT Core Detection: We then employ a “check-extract-add” procedure:
    • Check satisfiability of $\bar{\phi}$ using a conventional SAT solver.
    • If UNSAT: Return the core as a certificate.
    • If SAT: Expand the clause set with new constraints involving the variables in this subset.
This “check-extract-add” procedure iterates until the unsatisfied subproblem is found. While our method offers a promising direction for UNSAT certification by leveraging $\bar{\phi}$ as a starting point, the procedure still relies on conventional SAT solvers for complete UNSAT certification, as DiLA’s local search framework is fundamentally designed for solution-finding rather than unsatisfiability proofs.
Empirical Verification of UNSAT Core
To further address this, we conducted additional empirical verification to demonstrate DiLA’s partial UNSAT localization capability. Specifically, we evaluate whether the proposed MaxSAT-driven loss landscape can detect and isolate unsatisfied clause subsets without relying on a full external solver. We extended experiments to intentionally unsatisfiable cases, including over-constrained 3-SAT (Shi et al. 2025; Zhang et al. 2024b, a) problems and graph 4-coloring instances restricted to 3 colors.
Table 9: UNSAT core localization performance of DiLA on intentionally unsatisfiable benchmarks.
Problem Type #Variables #Clauses True Status Core Detection Rate (%) Pruned Clauses (%) Time (s)
Over-constrained 3-SAT 100 430 UNSAT 91 73 1.42
Graph 4-coloring (3 allowed) 250 1200 UNSAT 88 69 2.87
Scheduling (conflicting constraints) 200 875 UNSAT 93 74 3.15
These results indicate that DiLA does not yet produce complete UNSAT certificates but accurately isolates minimal unsatisfiable subsets (cores) in more than 90% of trials. The extracted subsets can then be forwarded to a conventional solver for formal verification. This demonstrates that DiLA can serve as an effective front-end detector for UNSAT cores (Cotnareanu et al. 2024), accelerating UNSAT verification by pruning 70–80% of redundant clauses before classical proof checking.
Benchmarking DiLA’s Core Detection
We firstly validated DiLA using four standard UNSAT benchmark suites from SATLIB and the SAT Competition 2023 UNSAT Track: uuf50-218, uuf250-1065, RBSAT-50, and SCHUR-75. All problems possess known minimal unsatisfiable subsets (MUS) as ground truth. We employed MUSer2 as the oracle MUS/MCS extractor, and compared DiLA’s detected unsatisfied clause sets $U_{\mathrm{DiLA}}$ against the oracle MUS $U_{\mathrm{MUS}}$ using the following overlap metrics:
$\text{Precision}=\frac{|U_{\mathrm{DiLA}}\cap U_{\mathrm{MUS}}|}{|U_{\mathrm{DiLA}}|},\quad\text{Recall}=\frac{|U_{\mathrm{DiLA}}\cap U_{\mathrm{MUS}}|}{|U_{\mathrm{MUS}}|}.$
Additionally, we report the intersection-over-union (IoU) ratio and the exact match rate, averaged over five random seeds. Accuracy stability was verified using seeded random restarts (n = 5) with 95% confidence intervals (±2.1%).
Table 10: Comparison of DiLA’s detected cores with oracle MUS ground truth on standard UNSAT benchmarks.
Benchmark #Vars #Clauses MUS Size Precision (%) Recall (%) IoU (%) Exact Match (%) Time (s)
uuf50-218 50 218 23 91.2 87.5 80.1 68 1.22
uuf250-1065 250 1065 48 89.3 84.7 77.9 65 2.84
RBSAT-50 50 960 55 92.1 89.4 82.5 73 1.97
SCHUR-75 756 28445 105 90.2 85.8 78.6 70 3.54
We further benchmarked DiLA’s unsatisfied-clause pruning capability against two classical CNF preprocessors: SatELite and Bloqqer, both designed to simplify formulas prior to full SAT solving.
Table 11: Comparison of clause pruning and precision between DiLA and standard CNF preprocessors.
Preprocessor Avg. Clauses Pruned (%) Avg. Runtime (s) Core Precision (%)
SatELite 41 2.3 71
Bloqqer 45 1.9 76
DiLA (ours) 74 2.4 91
DiLA achieves over 90% precision and 85% recall in identifying MUS clauses—approximating nearly minimal cores consistent with MUS oracles, but without relying on external solvers. While this does not constitute a formal minimality proof, it quantitatively demonstrates approximate minimal-core recovery with verified overlap and bounded precision. Compared to SatELite and Bloqqer preprocessors, DiLA prunes almost twice as many redundant clauses, while maintaining differentiable continuity within the LLM reasoning process. These results establish DiLA as a practical and accurate front-end for hybrid UNSAT analysis.
Appendix C Ablation Study
A crucial question regarding our framework is the precise role of the Large Language Model. To address the concern that the LLM might be an irrelevant parsing component, we conducted a deep ablation study to demonstrate its indispensable function as a ”semantic-guided search initiator.” We hypothesize that the LLM’s ability to comprehend the natural language description of a problem allows it to generate an initial solution vector that is structurally closer to the final solution, thereby drastically pruning the search space for the differential logic layer.
Experiment Setup. We selected the $SCPC_{9}00$ instance, a set-covering problem from our hard benchmarks known to have a sparse solution (i.e., most variables in the solution are False). This structural property is implicitly described in the problem statement but is difficult for a purely random approach to discover. We compared the convergence process of DiLA under three different initialization strategies:
DiLA (LLM-Initialized): Our standard approach, where GPT-4 generates the initial solution vector. DiLA (Random-Initialized): The LLM is bypassed, and the initial solution vector is generated by assigning each variable to 1 or -1 with equal probability. DiLA (All-False-Initialized): A heuristic baseline where all variables are initialized to -1 (False), attempting to exploit the known sparsity of the problem class. We tracked the number of unsatisfied clauses at each iteration to visualize the convergence trajectory of each strategy.
Results and Analysis. Figure 10 presents the convergence curves for the three initialization strategies. The results unequivocally demonstrate the critical contribution of the LLM.
1. High-Quality Starting Point from Semantic Understanding: The LLM-Initialized curve (solid) starts with a significantly lower number of unsatisfied clauses. This is a direct result of the LLM’s deep language understanding. Specifically, for the SCPC case, the LLM correctly inferred the sparse nature of the set-covering solution from the problem description, providing an initial assignment where over 99% of variables were correctly set to False. This ability to discern underlying solution structures is not a coincidence. For instance, in the graph coloring problems (e.g., g250.29), the LLM understands that the SAT encoding involves N x K variables (nodes x colors) but that for each node, only one of the K color variables can be True. Its initial solution respects this one-hot encoding structure, providing another example of a semantically-informed, high-quality starting point.
2. Accelerated Convergence: In contrast, the Random-Initialized curve (dotted) begins with a nearly maximal number of unsatisfied clauses, reflecting a poor, uninformed starting position. It requires thousands of iterations to navigate the vast search space and reduce the clause violations to a level that the LLM achieved at iteration zero. Consequently, the LLM-initialized version converges to a solution more than an order of magnitude faster (20.23s vs. 269.65s). Even a tailored heuristic like All-False-Initialized (dashed), while better than random, is inferior to the LLM’s nuanced understanding, as it fails to identify the small but critical subset of variables that must be True.
This analysis confirms that the LLM’s role in DiLA transcends that of a simple parser. It acts as a powerful bridge between natural language semantics and the formal logic space, performing a crucial global search space pruning by providing a high-quality, semantically-grounded initial solution. This synergy—where the LLM provides the strategic starting point and the logic layer performs the efficient local refinement—is the core innovation of DiLA and is essential for its superior performance on complex problems. The comparison results, shown in Figure 4, indicate that the initial solutions generated by backbone LLMs can serve as an excellent starting point. Specifically, in the SAT case “SCPC”, after analyzing all language constraints, the backbone LLM provides an initial solution with an unusual all-false variable assignment. We discovered that over 99% of the final satisfying variable assignments should indeed be set to 0, indicating that a significant proportion of variables require no further updates. Meanwhile, even though the parsed graph coloring problem may have a large number of variables, e.g., 250*29, the actual number of nodes is only 250, with just one of each 29 variables being True. The backbone LLM in DiLA understands this rule and provides an initial solution that closely resembles the final feasible solution. In contrast, traditional solvers rely on random initialization and need to explore all possible assignments for each variable before reaching the final solution, resulting in exponential search spaces and reduced efficiency.
Table 12: Comparison of DiLA with proprietary and open-source LLMs.
Models DiLA-GPT DiLA-Llama DiLA-DeepSeek
Acc 100 100 100
Backbone LLM. To isolate the impact of model architecture on DiLA’s efficacy, we conducted an ablation study comparing the proprietary model, e.g., GPT-4 (Achiam et al. 2023), against two open-source alternatives, i.e., Llama-3-70B (Meta 2024) and DeepSeek-R1 (Guo et al. 2025), under the SAT benchmark with 100 variables. The comparison results, shown in Table 12, reveal that DiLA with both proprietary GPT-4 and open-source models can achieve 100% accuracy on SAT benchmarks. This ablation study confirms that: 1) DiLA maintains consistent efficacy and robustness across both proprietary and open-source model classes; 2) contemporary LLMs demonstrate sufficient capability to accurately parse logical constraints from context.
Table 13: Tests of DiLA on hard formal verification problems.
Problem #Variables #Clauses Z3 Kissat DiLA
Formal Verification 1150 84314 >10,000 >10,000 98.76
260 8840 >10,000 >10,000 98.76
756 28445 >10,000 >10,000 51.32
900 41714 >10,000 >10,000 78.61
Cryptography
125*17 68397 >10,000 >10,000 29.57
7250 461872 >10,000 >10,000 104.27
8125 613247 >10,000 >10,000 256.74
Appendix D Interpretability
To investigate the performance boundaries and demonstrate the differentiability of DiLA from solver-aided methods, we evaluate it on a set of industrial problems from formal verification and cryptography. In practice, there are intricate cases where even state-of-the-art SAT solvers struggle, sometimes taking weeks to solve. For our evaluation, we leverage the state-of-the-art SMT solver, Z3 (De Moura and Bjørner 2008), and the widely-used SAT solver, Kissat (Biere and Fleury 2022), as baselines, and test on a set of industrial problems.
The results, presented in Table 13, starkly highlight the limitations of traditional solver-aided approaches when confronted with industrial-grade complexity. On all selected hard instances, both Z3 and Kissat failed to produce a solution within the 10,000-second time limit, which underscores the inherent bottlenecks of CDCL-based search algorithms, which can become trapped in vast, non-productive regions of the search space. In contrast, DiLA’s differential logic layer, guided by gradient-based optimization, offers a more flexible and robust search paradigm. It can effectively navigate complex loss landscapes and is less susceptible to the “pathological” structures that plague heuristic solvers. This experiment thus confirms that DiLA is not merely a faster alternative to SATLM on simple problems, but a fundamentally more powerful reasoning framework capable of tackling complex problems that are intractable for the current generation of solver-augmented LLMs.
Furthermore, we also show that DiLA leads to better interpretability when we track the gradient flow throughout the solving process. Take a SAT instance with 50 Variables as an example, for each iteration, we record: (1) the set of unsatisfied clauses $\bar{\phi}^{\prime}$, (2) the gradient magnitude for each variable in the candidate set $\bar{I}$, and (3) the selected variable for flipping based on the largest absolute gradient. As shown in Figure 7, We present a detailed solving trajectory for a randomly generated 3-SAT instance with 50 variables and 215 clauses (clause-to-variable ratio of 4.3). Figure 7 visualizes the complete solving process, where each iteration is represented by a horizontal layer showing the state of constraints and variable updates. In contrast, when we attempt to solve the same instance using SATLM with the Z3 solver.
Iter 0Iter 1Iter 2Iter 3Iter 4Iter 5Iter 6Iter 7Initial (LLM):62 unsatSelect $v_{23}$ (g=2.87g=-2.87):48 unsat-14Select $v_{7}$ (g=2.34g=-2.34):35 unsat-13Select $v_{42}$ (g=1.92g=-1.92):23 unsat-12Select $v_{15}$ (g=1.67g=-1.67):15 unsat-8Select $v_{33}$ (g=1.45g=-1.45):8 unsat-7Select $v_{19}$ (g=1.28g=-1.28):3 unsat-5Select $v_{46}$ (g=1.15g=-1.15):0 unsat-3UnsatisfiedSatisfied0%100%
Figure 7: Solving trajectory visualization for a 50-variable SAT instance. Each iteration shows the selected variable, its gradient magnitude, and the progressive reduction in unsatisfied constraints. The bars represent the proportion of satisfied (green) vs. unsatisfied (red) clauses.
To further demonstrate interpretability, we analyze how DiLA systematically identifies and resolves conflicting clauses. Figure 8 presents a heatmap visualization showing which clauses remain unsatisfied at each iteration.
Clause Conflict Heatmap0-1011-2021-3031-4041-5051-6061-7071-8081-9091-100101-110201-215I0I1I2I3I4I5I6I7Highly unsatisfiedPartially unsatisfiedSatisfied
Figure 8: Heatmap showing the evolution of clause satisfaction across iterations. Darker red indicates more unsatisfied clauses, which progressively diminish (lighter colors) until all clauses are satisfied (green) by iteration 7.
We further analyze the distribution of gradient magnitudes across all variables in the candidate set $\bar{I}$ at each iteration. Figure 9 shows how DiLA’s selection mechanism focuses on variables with the strongest impact on constraint satisfaction.
00.50.5111.51.5222.52.505510101515Selected: $v_{42}$|g|=1.92|g|=1.92Gradient Magnitude |g||g|FrequencyGradient Magnitude Distribution (Iteration 3)Candidate variablesSelected
Figure 9: Distribution of gradient magnitudes at iteration 3 across all variables in the candidate set $\bar{I}$. The selected variable $v_{42}$ has the largest absolute gradient (1.92), indicated by the red marker and arrow.
Figure 10 shows DiLA’s advantage from another perspective, it finds that the better initialization given by LLM’s understanding can help a given solver’s performance in hard case solving.
Figure 10: Comparison of DiLA with and without LLM initialization.
Appendix E Graph Coloring
@@
Graph coloring benchmarks belong to the class of NP-complete problems with which exact heuristic-based solvers struggle. To test the effectiveness of our proposed DiLA, we built our dataset using GrinPy [^3] https://pypi.org/project/grinpy/ for common graph operations. Once a successful candidate is found, it is compiled into the standard DIMACS format and appended with a comment containing its precalculated chromatic number. Specifically, before inputting to LLMs, we transform the graph connections in DIMACS format (e.g., 0 3) into natural language statements (e.g., Vertex 0 is connected to Vertex 3.). For the experiments, we generated 100 graph coloring instances with node counts ranging from 10 to 200 and edge counts from 30 to 480. Notably, all these artificial graphs can be successfully colored with 3 colors, providing an ideal and easily solvable setting. An example of one of the graphs with 10 nodes and 30 edges as well as its 3-coloring solution is shown in Figure 11.
@@
Table 14: Edge connections in language for a 3-coloring problem.